Show all work
Question 6 3 pts Using Case Study D, which of the following is the best way to organize the data for the NPZ-8 scores with a class width of 3? 04:59: 10-14 0-2; 3-5; 6-8; 9-11; 12-14
20Frequency%20Distributions%20&%20Histograms%20 Tables.pdf TABLE 2.10 Nurse Sick Nurse Sick Nurse Sick Number Days Number Days Number Days 15 16 17 18 19 7 20 21 14 TABLE 2.11 App. 16 18 App. m 220 225 f135 137 f 180 201 m 210 205 f 145 144 f 131 133 m 177 180 m 135 135 m 183 180 m 165 166 f 160 166 m 178 180 m 155 152 r 155 154 f 130 128 165 167 m 215 210 17 190 186 19 158 160 f 165 160 1 f 120 118 2 m 173 170 23 188 195 180 195 135 137 155 135 190 195 187 185 154 156 m 2 24 10 m 26 12 29 30 m 15
Answer #1

min = 0
max=14
class width = 3
(14-0)/3 = 14/3 = 4.666 hence 5 classes

option C) 0-2,3-5,6-8,9-11,12-14 is correct